6. Show that xe−x


2


is an eigenfunction of the linear operator d


2


dx2 − 4x


2


.


What is the eigenvalue?


1
Expert's answer
2021-12-27T13:03:30-0500

We have the linear operator "\hat A=\frac{d^2}{dx^2}-4x^2" .

"\psi (x)=xe^{-x^2}" is an eigenfunction of "\hat A" , if "\hat A\psi =\lambda \psi" .


"\hat A\psi=\frac{d^2}{dx^2}\big(xe^{-x^2}\big)-4x^2\cdot xe^{-x^2}=\frac{d}{dx}\big( e^{-x^2}-2x^2e^{-x^2}\big)-4x^3e^{-x^2}= -2xe^{-x^2}-4xe^{-x^2}+4x^3e^{-x^2}-4x^3e^{-x^2}=-6xe^{-x^2}=-6\psi"


So, "\psi(x)" is an eigenfunction of "\hat A" and the eigenvalue is "\lambda=-6" .